OpenStudy (anonymous):

If = 63° and Z = 19 cm, what is the value of X to the nearest tenth of a centimeter? A. 16.9 cm B. 41.9 cm C. 8.6 cm D. 8 cm

OpenStudy (anonymous):

picture below

OpenStudy (cwrw238):

which angle is 63° ?

OpenStudy (anonymous):

acute, its at the top u see it in the picture?

OpenStudy (cwrw238):

ok use trigonometry : x is adjacent to the angle of 63 degrees and z is the hypotenuse of the right angled triangle so cos 63 = adjacent / hypotenuse = x / 19 cos 63 = x / 19 x = 19 * cos 63

OpenStudy (cwrw238):

now its a calculator job to finish this off if the diagram is roughly to scale it looks like x = either 8 or 8.6

OpenStudy (anonymous):

ok thanks again!